Problem #1

Five jobs are to be processed through a single machine. The processing times and due dates are given here. Assume all released dates are zero.

Job

1

2

3

4

5

Processing time

3

6

5

4

2

Due date

4

8

12

21

15

In each of the following cases determine the sequence in which she should perform and compute Mean Flow Time, Average Tardiness and the number of tardy jobs.

A) Shortest Processing Time

B) Early Due Date
Answer #1

A)

SPT

Job Processing time Due date Start Finish Flow time Tardiness
5 2 15 0 2 2 0
1 3 4 2 5 5 1
4 4 21 5 9 9 0
3 5 12 9 14 14 2
2 6 8 14 20 20 12
Avg. 10 3

Average flow time = 10

Average tardiness = 3

Average number of tardy jobs = 3

B)

EDD

Job Processing time Due date Start Finish Flow time Tardiness
1 3 4 0 3 3 0
2 6 8 3 9 9 1
3 5 12 9 14 14 2
5 2 15 14 16 16 1
4 4 21 16 20 20 0
Avg. 12.4 0.8

Average flow time = 12.4

Average tardiness = 0.80

Average number of tardy jobs = 3

Calculations:

1 SPT Processing Due Job Start Finish Flow time Tardiness date time 2 3 5 ЕMAX(0, EЗ-С3) 15 =D3+B3 ЕВЗ =MAX(0,E4-C4) 4 1 3 4